OpenStudy (anonymous):

2( log 2 x − log y ) − ( log 3 + 2 log 5)

OpenStudy (anonymous):

what are you supposed to do about it?

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

it is not an equation. you cant solve it.

OpenStudy (anonymous):

cant it be simplified?

OpenStudy (anonymous):

You want to simplify it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it log (2x), or log x to base 2?

OpenStudy (anonymous):

it is log((2x/y)^2)-log 75

OpenStudy (anonymous):

log 2x

OpenStudy (anonymous):

not that log a - log b = log(a/b) and log a +log b = log(ab)

OpenStudy (anonymous):

how did u get the answer dhatraditya?

OpenStudy (anonymous):

note*

OpenStudy (anonymous):

note also that xlog a = log a^x

OpenStudy (anonymous):

\[2 (\log(2x)-\log y)-(\log 3+2\log 5)=2(\log {2x \over y})-\log(3*5^2)\] =\[\log({2x \over y})^2-\log(75)=\log({4x^2 \over 75y^2})\]

OpenStudy (anonymous):

I think this is the simplest it can be.

OpenStudy (anonymous):

so its simplified form is log ((2x/y)^2/75))

OpenStudy (anonymous):

What AnwarA has.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

how did u get the 75?

OpenStudy (anonymous):

3*5^2=3(25)=75

OpenStudy (anonymous):

as I explained, x log a = log a^x so 2 log 5 = log 5^2 = log 25 log a + log b = log (ab) now, log 3+ log 25 = log (3 times25) = log 75

OpenStudy (anonymous):

i understand now.thx

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

solve log base x (5x+24)=2

OpenStudy (anonymous):

x = 8 or -3

OpenStudy (anonymous):

since logs are not defined for negative numbers, x = 8

OpenStudy (anonymous):

do you know how I arrived at the answer, leona?

OpenStudy (anonymous):

?how did u work it out

OpenStudy (anonymous):

do you know how logarithms work?

OpenStudy (anonymous):

i know some of the rules

OpenStudy (anonymous):

tell me what you know about logs? what do you understand by logs?

OpenStudy (anonymous):

i know that the logarithm of 1 regardless of the base is one

OpenStudy (anonymous):

if i told you log 32 to the base 4 is 2.5, what do I mean by that?

OpenStudy (anonymous):

that 4 ^2.5 is 32

OpenStudy (anonymous):

exactly! good! have you tried that in your calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you understand that any number can be represented as a power of another number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay good. so what do i mean when i say log(5x+24) to the base x is 2?

OpenStudy (anonymous):

it means that x raise to the power 2 equal (5x+24)

OpenStudy (anonymous):

I should have said any POSITIVE number can be represented as apower of any other number (the base)

OpenStudy (anonymous):

right, so x^2 = 5x+24

OpenStudy (anonymous):

this is a simple quadratic equation

OpenStudy (anonymous):

should i square the x^2?

OpenStudy (anonymous):

why would you want to do that?

OpenStudy (anonymous):

i dont know how to remove the x

OpenStudy (anonymous):

okay have you solved quadratic equations before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you see that x^2 = 5x+24 can be rearranged so that it is in the form ax^2+bx+c = 0?

OpenStudy (anonymous):

ohh yea

OpenStudy (anonymous):

after that it is straightforward factoring. Remember that you will get a negative root. You cannot use the negative root, because as I explained before, logarithms for negative bases are generally discarded.

OpenStudy (anonymous):

what r the two factors? cant find them

OpenStudy (anonymous):

post your equation.

OpenStudy (anonymous):

x^2-5x-24=0

OpenStudy (anonymous):

i got 8

OpenStudy (anonymous):

yeah!!!!!!

OpenStudy (anonymous):

did you use the quadratic formula or the factoring method?

OpenStudy (anonymous):

quadratic

OpenStudy (anonymous):

okay. anything that works, although in this case factoring is more elegant and efficient. remember, you have to discard the x= -3 root

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

okay, do you know the reason why you discarded -3?

OpenStudy (anonymous):

because log not defined for negative numbers

OpenStudy (anonymous):

no, its because logs with negative bases only work for even powers. anyway, good job! have a medal.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

do a little homework assignment for me. Google "why logarithms are not defined for negative bases" and "why logarithms are not defined for negative numbers". it will help you understand logs much better. all the best!

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

ill do that

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

are u going to be online in the night?

OpenStudy (anonymous):

i dont know. I generally come here in my free time to help out.

OpenStudy (anonymous):

ohh cause i ave exam 2marow and i need more help with other topics

OpenStudy (anonymous):

oh.. you can post your questions here so that I get notifications and I'll try to answer them before tomorrow.

OpenStudy (anonymous):

ok am going to try this question log base x (5x+14)=2

OpenStudy (anonymous):

okay, lemme know how it works out. you should get 7

OpenStudy (anonymous):

logarithm are not defined for negative values because the log and In graphs have a vertical asymptotic line at x=0 which is the lower boundary of the graph ..in fact the limit as x approaches zero from the right hand side ,the function approaches infinity and that makes the zero point the maximum tolerable value for the x values less than zero

OpenStudy (anonymous):

i got 7....using the factoring method

OpenStudy (anonymous):

okay good.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your latest logarithm question will go like : \[5x +14=x^{2}\]

OpenStudy (anonymous):

so \[x ^{2}-5x-14=0 and then you solve with the quadratic formula \]

OpenStudy (anonymous):

need help with this 1....log base 4 (1) +log (25)

OpenStudy (anonymous):

so x= 7 or x=-2 but when you substitute the x=-2 the equation doesnt make sence so x=7 is the answer hope that was helpful

OpenStudy (anonymous):

okay, what does log base 4 (1) mean?

OpenStudy (anonymous):

log to base 4 to a number = 1

OpenStudy (anonymous):

log base 4 of 1 means what power will you raise the number 4 to have a value of 1 so its like : 1=4\[^{x}\]

OpenStudy (anonymous):

it means that 4 raised to a number should be equal to 1. what is that number? we know that any number raised to the power 0 is 1. so log base 4 (1) is 0. so your equation simplifies to just log 25

OpenStudy (anonymous):

like 4\[4^{x}=1 \]

OpenStudy (anonymous):

that means x=0

OpenStudy (anonymous):

because

OpenStudy (anonymous):

\[4^{0}\]

OpenStudy (anonymous):

ohh yeah thats a rule of logs

OpenStudy (anonymous):

yes. you should remember the basic rules of logs. Also, understand what logarithms mean. That way, even if you get stuck in your test, you can work it out.

OpenStudy (anonymous):

so 5^2=25

OpenStudy (anonymous):

right, log 25 = log 5^2 = ?

OpenStudy (anonymous):

yeah so the final answer is log 5^2=25

OpenStudy (anonymous):

log 25 without a base specified means the base is a standard base of ten (10)

OpenStudy (anonymous):

so are you saying that 10^25 = 5^2, leona?